# Level-set shape optimization

### v1

@@ 1 Introduction
Polytopic meshes also greatly simplify fitted methods. In [35, 36], the authors discretize the level-set function using standard finite elements on hexahedral meshes, truncate the hexahedra with the zero-level set to generate a fitted polytopic mesh, employ modified trilinear and moving-least squares basis functions to compute the state, and solve the level-set equation using “essentially non-oscillatory” (ENO) schemes. In [25], the authors present a level-set shape optimization based on the topological derivative and generate fitted polytopic meshes by intersecting an initial Voronoi tessellation with the zero-level set. The state equation is solved using the symmetric Interior Penalty discontinuous Galerkin (IPdG) method on the polytopic mesh. The level-set equation is regularized by adding a small regularizing amount of diffusion and is solved on the polygonal Voronoi mesh using a symmetric IPdG method. This approach does not include the equivalent of the shape gradient equation because the topological derivative is applied pointwise.
@@
These works demonstrated some of the advantages polytopic meshes offer in the context of shape optimization such as higher element anysotropy or facilitated local refinement. However, polytopic methods have much more to offer. The last ten years or so have seen the development of a rich theory (suported by efficient implementations) of discontinuous Galerkin (dG) methods on general, possibly curved, polytopic elements [14, 12, 13, 19, 11]. In particular, the element-wise discontinuous nature of the approximations spaces permitted by dG methods allows for extremely general element shapes. In this work, we build on this enhanced geometric flexibility. Specifically, we propose a unified polytopic discontinuous Galerkin framework for fitted level-set shape optimization based on agglomeration of elements of a fine underlying simplicial mesh. The unified approach allows decoupling the discretization of the level-set equation, the discretization of the shape gradient equation, and the discretization of the state (and adjoint) equations. This approach streamlines and facilitates the construction of coarse agglomerated polytopic elements that fit the zero-level set to great accuracy and that, in combination with higher-order dG methods, allow the rapid solution of the shape gradient and of the state equations. Transferring information between a coarse polytopic mesh and the underlying fine simplicial mesh is also straightforward. This enables solving the level-set equation on the fine underlying simplicial mesh using standard, explicit, and stable Runge-Kutta dG (RKdG) methods, to a computational cost comparable to that of solving the shape gradient equation. We note that in [34] the authors evidence that using RKdG methods for the level-set equation in divergence form is very efficient, accurate, stable. Additionally, in contrast to Hamilton-Jacobi ENO methods, RKdG methods do not require non-standard procedures such as reinitialization. In this work, we apply RKdG methods directly to the level-set equation in non-divergence form, without any observed detrimental effect to the stability of the method.
@@
The remainder of this work is organized as follows. In Section 2, we briefly review the level-set shape optimization method. In Section 3, we discuss the discontinuous Galerkin approximation of the shape gradient, the state, and the level-set equation, paying extra care paid to the consistent evaluation of the shape derivative on discontinous test functions. In Section 4, we present a selection of numerical experiments that demonstrate the efficacy of the new method. We consider both unconstrained and PDE-constrained shape optimization test cases and use a steepest-descent optimization algorithm with backtracking and Armijo stopping rule. The method is implemented in Python and the code is available on the Zenodo archive at [24]. Finally, in Section 5 we draw conclusions and outline future research directions.
@@ 3 Numerical level-set shape optimization using dG methods
The numerical realization of the level-set shape optimization method Eq. 2 requires, at the very least, the discretization of the level set equation Eq. 1 and of the shape gradient equation Eq. 3. A very popular approach is to employ finite differences to solve Eq. 1 in combination with a standard conforming Galerkin method to solve Eq. 3. Instead, here we develop a unified polytopic discontinuous Galerkin method framework. The so-called polytopic (i.e., polygonal for $d=2$ or polyhedral for $d=3$) discontinuous Galerkin methods have been developed to be able to operate on meshes comprising essentially arbitrary element shapes. In [14, 12, 13, 11], it has been shown that it is possible to construct stable discontinuous Galerkin methods on such extreme mesh scenarios. A pertinent feature of these methods is that the cardinality of the local elemental degrees of freedom are independent of the element shape, that is element geometry and number of faces/edges/vertices. This feature is naturally very attractive for shape optimization frameworks, whereby it is extremely important to strike a balance between total computational complexity, by using as coarse mesh as possible, and geometric feature resolution. As we shall see below, by allowing meshes with elements with essentially arbitrary shapes, we can follow zero-level sets with extremely high fidelity without requiring highly locally refined meshes in the zero-level set neighborhoods.
@@
$H^{1}(\mathscr{T})\coloneqq\{v\in L^{1}(D,\mathbb{R}^{d})\,:v|_{T}\in H^{1}(T,\mathbb{R}^{d})\text{ for all }T\in\mathscr{T}\},$
@@ 3.2 The discontinuity-penalization parameter
For completeness, we now give a precise formula for the discontinuity-penalization function $\sigma$ appearing in the interior penalty discontinuous Galerkin (IP-dG) formulations above. The stability and error analysis of the IP-dG method under this choice of penalization, as well as a detailed discussion on the practical relevant of this choice, can be found in [13, 11].
@@ 3.4 Runge-Kutta discontinuous Galerkin discretization of the level-set equation
We note that, employing a spatial discretization of Eq. 1 on $\tilde{\mathscr{T}}$ (or a refinement thereof!), is not necessarily detrimental to the overall complexity of the method. This is due to the explicit timestepping nature of RKdG, since no global linear system is required to be solved on each time-step, and since quadratures take place on simplicial subdivisions. On the contrary, it is advantageous in terms of accuracy to discretize the level-set equation Eq. 1 over a finer mesh than the shape gradient equation to be able to fine-tune the evolution of the shape.
@@ 3.5 Polytopic dG discretization of the level-set shape optimization method
Figure 2: Example of local refinement method in 2D. Upon each edge that is intersected by the zero-level set $\{\phi=0\}$ (red curved line), we interpolate the level set function and determine its zero, where we to append new node (blue dots) used to split the triangle in (at most) three subtriangles.
@@ 4.1 Optimization algorithm
for a chosen Hilbert space ${\cal H}$. In the following numerical experiments, we use ${\cal H}=H^{1}_{0}(D;\mathbb{R}^{d})$, but other choices are possible. The final time $\epsilon_{k}$ must be chosen such that $\varphi^{k}(t)$ exists for every $t\in(0,\epsilon_{k})$. Note that both $\epsilon_{k}$ and the time step $\Delta t_{k}$ used to approximate $\varphi^{k}$ depend on the $W^{1,\infty}(D;\mathbb{R}^{d})$-norm of $\nabla J(\Omega_{\phi^{k}})$, which possibly changes at each iteration $k$. In the following numerical experiments, we select $\Delta t_{k}$ to satisfy the CFL-condition and perform at most $M\in\mathbb{N}$ time steps, thus implicitly setting $\epsilon_{k}=M\Delta t_{k}$. The stopping time $t^{k}$ is selected in a backtracking fashion as the largest in the set $\{m\Delta t_{k}\,:m\in\mathbb{N},0<m\leq M\}$ such that the Armijo condition
@@
Figure 3: Workflow of the agglomerated polytopic dG level-set shape optimization method. A base triangular mesh is locally refined to better fit the zero-level set $\phi=0$. The resulting mesh is subsequently agglomerated to compute an approximate solution to the shape gradient equation Eq. 26. This solution is then interpolated on the locally refined triangular mesh, where the level-set equation Eq. 25 is solved using a time-stepping scheme. When the time-stepping scheme stops, the locally refined mesh is adjusted to better fit the zero-level set of the updated level-set function $\phi$. This procedure is repeated until the convergence or stopping criteria are met.
Figure 4: Extension of the agglomerated polytopic dG level-set shape optimization workflow presented in Fig. 3 to PDE-constrained shape optimization problems. To compute the shape gradient $\nabla J$, we first extract the state equation’s computational domain and compute the state solution and its lift $R(u)$.
@@ 4.2 Unconstrained test case
To solve the level-set equation Eq. 25, we partition the hold-all domain $D$ into 1800 triangles, discretize in space using RKdG with spatial basis functions of degree $p=2$ and timestepping via Heun’s method with time step $\Delta t_{k}=1/2600$, with at most $M=30$ time steps per iteration. Choosing a very small time step $\Delta t_{k}$ is not a requirement of the dG method employed. This choice is simply due to the desire to generate “smooth” convergence histories for plotting purposes and to provide a “smallest step” stopping criteria for the backtracking algorithm. To solve the shape gradient equation Eq. 26, we agglomerate the locally refined triangular mesh (cf. Fig. 3) into 200 polytopic elements and employ polynomials of degree $p=2$. The computation of the shape derivate $dJ$ in the discontinuous Galerkin framework is discussed in Example 3.2, see Eq. 17. To assemble the various Galerkin matrices, we employ quadrature rules of order four both on triangles and on edges, which compute the involved integrals exactly (up to round-off errors).
@@
Some shape iterates, starting from a disc of radius 0.51 centered at the origin, are displayed in Fig. 5. We observe that the algorithm retrieves a good approximation of the optimal (and not connected) shape. To assess the advantage of the agglomerated polytopic dG approach to compute shape derivatives, we compare it with a more classical dG approach based on triangular meshes. For the latter, we also employ an adaptive quadrature based on locally refined meshes to evaluate the shape derivative formula $dJ$ accurately on triangles that intersect the level-set boundary $\phi=0$. For both methods, we solve the level-set equation Eq. 25 on a fine underlying simplicial mesh as in the previous numerical experiment. Figure 6 summarizes the qualitative and quantitative comparison between the two approaches when using 50 and 200 triangular or polytopic elements. We observe that the adaptive nature of the agglomerated polytopic approach allows it to achieve accurate results also with coarser meshes.
@@
Figure 5: Evolution of shapes for the unconstrained test case. The initial domain, a disc, elongates and eventually splits into two domains before converging to the optimal configuration. The contour plots represent the normalized contours of the Euclidean norm of the gradient $\nabla J$. Note that the polytopic mesh changes at each iteration.
@@
Figure 6: Comparison between the agglomerated polytopic dG approach and a more classical approach based on triangular elements with adaptive quadrature. The figure on the left depicts the optimized shape, the optimal shape, and the normalized contours of the Euclidean norm of the gradient $\nabla J$. The figure on the right denotes the evolution of the objective function (with markers denoting gradient evaluations) versus the pseudo-time $t$. We observe that the agglomerated polytopic dG approach with 50 elements perform qualitatively better than (and is quantitatively comparable to) the adaptive quadrature approach with 200 elements.

### v2

@@ 1 Introduction
Polytopic meshes also greatly simplify fitted methods. In [37, 38], the authors discretize the level-set function using standard finite elements on hexahedral meshes, truncate the hexahedra with the zero-level set to generate a fitted polytopic mesh, employ modified trilinear and moving-least squares basis functions to compute the state, and solve the level-set equation using “essentially non-oscillatory” (ENO) schemes. In [27], the authors present a level-set shape optimization based on the topological derivative and generate fitted polytopic meshes by intersecting an initial Voronoi tessellation with the zero-level set. The state equation is solved using the symmetric Interior Penalty discontinuous Galerkin (IPdG) method on the polytopic mesh. The level-set equation is regularized by adding a small regularizing amount of diffusion and is solved on the polygonal Voronoi mesh using a symmetric IPdG method. This approach does not include the equivalent of the shape gradient equation because the topological derivative is applied pointwise. A different approach is presented in [17], where the authors combine polytopic meshes induced by a Laguerre diagram with the virtual element method and realized a shape and topology optimization method by acting directly on the Laguerre diagram parameters.
These works demonstrated some of the advantages polytopic meshes offer in the context of shape optimization such as higher element anisotropy or facilitated local refinement. However, polytopic methods have much more to offer. The last ten years or so have seen the development of a rich theory (supported by efficient implementations) of discontinuous Galerkin (dG) methods on general, possibly curved, polytopic elements [14, 12, 13, 21, 11]. In particular, the element-wise discontinuous nature of the approximations spaces permitted by dG methods allows for extremely general element shapes. The ability to use elements of essentially arbitrary shape with the same number of local numerical degrees of freedom as those used on simplicial elements [14, 12, 13] effectively decouples the element shape from its approximation capabilities. This is extremely attractive in the context of shape optimization, since coarse polytopic elements can accurately represent complex shapes, in contrast to what is possible with coarse simplicial meshes. Moreover, hanging nodes are naturally embedded in the polytopic mesh paradigm.
@@
In this work, we build on this enhanced geometric flexibility. Specifically, we propose a polytopic discontinuous Galerkin framework for fitted level-set shape optimization based on agglomeration of elements of a fine underlying simplicial mesh. The approach allows decoupling the discretization of the level-set equation, the discretization of the shape gradient equation, and the discretization of the state (and adjoint) equations. This approach streamlines and facilitates the construction of coarse agglomerated polytopic elements that fit the zero-level set to great accuracy and that, in combination with higher-order dG methods, allow the rapid solution of the shape gradient and of the state equations. Transferring information between a coarse polytopic mesh and the underlying fine simplicial mesh is also straightforward. This enables solving the level-set equation on the fine underlying simplicial mesh using standard, explicit, and stable Runge-Kutta dG (RKdG) methods, to a computational cost comparable to that of solving the shape gradient equation. We note that in [36] the authors evidence that using RKdG methods for the level-set equation in divergence form is very efficient, accurate, stable. Additionally, in contrast to Hamilton-Jacobi ENO methods, RKdG methods do not require non-standard procedures such as reinitialization. In this work, we apply RKdG methods directly to the level-set equation in non-divergence form, without any observed detrimental effect to the stability of the method.
@@
The remainder of this work is organized as follows. In Section 2, we briefly review the level-set shape optimization method. In Section 3, we discuss the discontinuous Galerkin approximation of the shape gradient, the state, and the level-set equation, paying extra care to the consistent evaluation of the shape derivative on discontinous test functions. In Section 4, we present a selection of numerical experiments that demonstrate the efficacy of the new method. We consider both unconstrained and PDE-constrained shape optimization test cases and use a steepest-descent optimization algorithm with backtracking and Armijo stopping rule. The method is implemented in Python and the code is available on the Zenodo archive at [26]. Finally, in Section 5 we draw conclusions and outline future research directions.
@@ 2 The level-set shape optimization method
Remark 2.2.
As described in [18], it is also possible to compute steepest descent directions in $W^{1,\infty}(D,\mathbb{R}^{d})$. These descent directions can be used to drive the evolution of the level-set function $\varphi^{k}$ in Eq. 2 instead of using the negative shape gradient $-\nabla J(\Omega_{\phi^{k}})$. To the best of our understanding, computing such descent directions may be computationally more involved than solving Eq. 3. Although we are not aware of their use in level-set based shape optimization, in other frameworks it has been observed that $W^{1,\infty}(D,\mathbb{R}^{d})$-steepest direction based methods can outperform Hilbert space-based counterparts, especially when optimizing domains with corners. Since a comparison between different choices of descent directions is not the focus of this work, for simplicity we compute descent directions by solving Eq. 3. We highlight that, as presented below, the shape gradient equation Eq. 3 can be solved efficiently using polytopic meshes due to their reduction of computational complexity.
@@ 3 Numerical level-set shape optimization using dG methods
The numerical realization of the level-set shape optimization method Eq. 2 requires, at the very least, the discretization of the level set equation Eq. 1 and of the shape gradient equation Eq. 3. A very popular approach is to employ finite differences to solve Eq. 1 in combination with a standard conforming Galerkin method to solve Eq. 3. Instead, here we develop a polytopic discontinuous Galerkin method framework. The so-called polytopic (i.e., polygonal for $d=2$ or polyhedral for $d=3$) discontinuous Galerkin methods have been developed to be able to operate on meshes comprising essentially arbitrary element shapes. In [14, 12, 13, 11], it has been shown that it is possible to construct stable discontinuous Galerkin methods on such extreme mesh scenarios. A pertinent feature of these methods is that the cardinality of the local elemental degrees of freedom are independent of the element shape, that is element geometry and number of faces/edges/vertices. This feature is naturally very attractive for shape optimization frameworks, whereby it is extremely important to strike a balance between total computational complexity, by using as coarse mesh as possible, and geometric feature resolution. As we shall see below, by allowing meshes with elements with essentially arbitrary shapes, we can follow zero-level sets with extremely high fidelity without requiring highly locally refined meshes in the zero-level set neighborhoods.
@@
${H^{1}(\mathscr{T})\coloneqq\{v\in L^{1}(D)\,:v|_{T}\in H^{1}(T)\text{ for all}T\in\mathscr{T}\},}$
@@ 3.2 The discontinuity-penalization parameter
For completeness, we now give a precise formula for the discontinuity-penalization function $\sigma$ appearing in the interior penalty discontinuous Galerkin (IP-dG) formulations above. The stability and error analysis of the IP-dG method under this choice of penalization, as well as a detailed discussion on the practical relevance of this choice, can be found in [13, 11].
@@ 3.4 Runge-Kutta discontinuous Galerkin discretization of the level-set equation
We note that, employing a spatial discretization of Eq. 1 on $\tilde{\mathscr{T}}$ (or a refinement thereof!), is not necessarily detrimental to the overall complexity of the method. This is due to the explicit time-stepping nature of RKdG, since no global linear system is required to be solved on each time-step, and since quadratures take place on simplicial subdivisions. Indeed, the main computational cost (associated with the time-stepping component of an explicit RKdG method) consists in assembling elemental mass matrices and solving the resulting block-diagonal linear systems. Using $L^{2}$-orthogonal bases, which are available for simplices, the mass matrices become diagonal, making this computational costs negligible. Of course, it is by all means possible to employ general polytopic meshes for the discretization of the level-set equation [12]. In this case, assembly of elemental mass matrices typically involves quadrature computations on a simplicial subdivision of the polytopic mesh, in which case the computational cost is the same as computing directly on the simplicial subdivision. For large problems, this can be significantly accelerated via GPU implementations [21]. Using polytopic meshes may lead to fewer but larger blocks compared to a discretization on the underlying simplicial mesh. However, the resulting polytopic elemental mass matrices are generally not diagonal, and so the fast solution aspect may be compromised. We also note that the spatial accuracy is generally driven by elements’ diameters, and so solutions computed on polytopic meshes may be less accurate than counterparts obtained on their simplicial subdivision. Therefore, given the comparable computational cost and potentially higher accuracy, we opt for solving the level-set equation on the simplicial subdivision $\tilde{\mathscr{T}}$.
@@ 3.5 Polytopic dG discretization of the level-set shape optimization method
Figure 2: Example of local refinement method in 2D. Upon each edge that is intersected by the zero-level set $\{\phi=0\}$ (red curved line), we interpolate the level set function and determine its zeros, where we append new nodes (black dots) used to split the triangle in (at most) three subtriangles. To obtain a more accurate representation of the zero-level set, one could first subdivide elements that intersect the zero-level set and then apply the local refinement method to newly generated elements that intersect the zero-level set, possibly in combination with a higher-order moving-mesh shape optimization method [42].
@@ 4.1 Optimization algorithm
for a chosen Hilbert space ${\cal H}$. In the following numerical experiments, we use ${\cal H}=H^{1}_{0}(D;\mathbb{R}^{d})$, but other choices are possible. The final time $\epsilon_{k}$ must be chosen such that $\varphi^{k}(t)$ exists for every $t\in(0,\epsilon_{k})$. Note that both $\epsilon_{k}$ and the time step $\Delta t_{k}$ used to approximate $\varphi^{k}$ depend on the $W^{1,\infty}(D;\mathbb{R}^{d})$-norm of $\nabla J(\Omega_{\phi^{k}})$, which possibly changes at each iteration $k$. In the following numerical experiments, we select $\Delta t_{k}$ to satisfy the CFL-condition and perform at most $M\in\mathbb{N}$ time steps, thus implicitly setting $\epsilon_{k}=M\Delta t_{k}$. The stopping time $t^{k}$ is selected in a backtracking fashion as the largest in the set $\{{5}m\Delta t_{k}\,:m\in\mathbb{N},0<{5}m\leq M\}$ such that the Armijo condition
@@
Figure 3: Workflow of the agglomerated polytopic dG level-set shape optimization method. A base triangular mesh (left) is locally refined to better fit the zero-level set $\phi=0$. The resulting mesh (middle) is subsequently agglomerated into a polytopic mesh (right) to compute an approximate solution to the shape gradient equation Eq. 26. The solution $\nabla J$ is then recovered (via interpolation and nodal averaging) on the locally refined triangular mesh, where the level-set equation Eq. 25 is solved using a time-stepping scheme. When the time-stepping scheme stops, the base mesh is locally refined mesh to better fit the zero-level set of the updated level-set function $\phi$. This procedure is repeated until the convergence or stopping criteria are met.
Figure 4: Extension of the agglomerated polytopic dG level-set shape optimization workflow presented in Fig. 3 to PDE-constrained shape optimization problems. The workflow begins as in Fig. 3. To compute the shape gradient $\nabla J$, we first extract the state equation’s computational domain and compute the state solution $u$. The state solution is then recovered (via interpolation of $u$, $\nabla u$, and the lift $R(u)$) on the polytopic mesh used to compute the shape gradient $\nabla J$. Then, the workflow continues as in Fig. 3.
@@ 4.2 Unconstrained test case
To solve the level-set equation Eq. 25, we partition the hold-all domain $D$ into 1800 triangles, discretize in space using RKdG with spatial basis functions of degree $p=2$ and timestepping via Heun’s method with time step $\Delta t_{k}=1/2600$, with at most $M={150}$ time steps per iteration. Choosing a very small time step $\Delta t_{k}$ is not a requirement of the dG method employed. This choice is simply due to the desire to generate “smooth” convergence histories for plotting purposes and to provide a “smallest step” stopping criteria for the backtracking algorithm. To solve the shape gradient equation Eq. 26, we agglomerate the locally refined triangular mesh (cf. Fig. 3) into 200 polytopic elements and employ polynomials of degree $p=2$. Table 1 reports a numerical validation of the polytopic discontinuous Galerkin approximation of the shape gradient. The computation of the shape derivate $dJ$ in the discontinuous Galerkin framework is discussed in Example 3.2, see Eq. 17. To assemble the various Galerkin matrices, we employ quadrature rules of order four both on triangles and on edges, which compute the involved integrals exactly (up to round-off errors).
Table 1: Numerical validation of the polytopic discontinous Galerkin approximation of the shape gradient $\nabla J$ in the unconstrained test case. The computational domain is a square of edge length 2 and the zero-level set is a circle of radius 0.52 (both centered at the origin, as depicted in white in the most left image in Fig. 5). In the table, the index $\mathrm{i}$ denotes the level of refinement, $\mathrm{N}$ denotes the number of polytopic elements, and $\mathrm{err}_{l}$ and $\mathrm{err}_{q}$ denote the $L^{2}$-error for affine and quadratic polygonal finite element approximations, respectively. The approximate reference solution has been computed using standard continuous quadratic finite elements on a simplicial mesh with 4944 elements. We observe that the approximate algebraic converge rates computed using the formula $\mathrm{rate}(i)=\log(\mathrm{err}(\mathrm{i})/\mathrm{err}(\mathrm{i}-1))/\log(\mathrm{N}(\mathrm{i})/\mathrm{N}(\mathrm{i}-1))$ are close to $-1$ and $-1.5$, which are the expected rates for finite element approximations based on affine and quadratic polynomials, respectively.
$\mathrm{i}$ $\mathrm{N}$ $\mathrm{err}_{l}$ $\mathrm{rate}_{l}$ $\mathrm{err}_{q}$ $\mathrm{rate}_{q}$
1 36 1.34e-3 - 3.09e-4 -
2 144 4.06e-4 -0.86 6.70e-5 -1.10
3 576 1.35e-4 -0.79 7.25e-6 -1.60
4 2304 3.86e-5 -0.91 1.04e-6 -1.40
@@
Some shape iterates, starting from a disc of radius 0.51 centered at the origin, are displayed in Fig. 5. The normalized pseudo-time evolution of the objective function is depicted in Fig. 7 We observe that the algorithm retrieves a good approximation of the optimal (and not connected) shape. To assess the advantage of the agglomerated polytopic dG approach to compute shape derivatives, we compare it with a more classical dG approach based on triangular meshes. For the latter, we also employ an adaptive quadrature based on locally refined meshes to evaluate the shape derivative formula $dJ$ accurately on triangles that intersect the level-set boundary $\phi=0$. For both methods, we solve the level-set equation Eq. 25 on a fine underlying simplicial mesh as in the previous numerical experiment. Figures 6 and 7 summarizes the qualitative and quantitative comparison between the two approaches when using 50, 200, and 450 triangular or polytopic elements. The quantitative comparison is performed both in terms of the objective functions $J$ and a numerical approximation of the distance between zero-level sets defined as $\max_{x\in\{\varphi=0\}}\min_{y\in\{f=0\}}|x-y|$. We observe that the adaptive nature of the agglomerated polytopic approach allows it to achieve accurate results also with coarser meshes.
Refer to captionRefer to caption
Refer to captionRefer to caption
$n=0$, $t=0$$n=9$, $t=0.52$$n=14$, $t=0.81$$n=41$, $t=1.45$
@@
Figure 5: Evolution of shapes for the unconstrained test case. The variables $n$ and $t$ denote the number of gradient evaluations and the pseudo-time, respectively. The initial domain, a disc (drawn in white), elongates and eventually splits into two domains before converging to the optimal configuration (drawn in red). The contour plots represent the normalized contours of the Euclidean norm of the gradient $\nabla J$. Note that the polytopic mesh changes at each iteration.
Figure 6: Qualitative comparison (based on the unconstrained test case) between the agglomerated polytopic dG approach and a more classical approach based on triangular elements with adaptive quadrature. The optimal solution is the union of two disjoint disk-shaped domains depicted in red. The optimized shapes are depicted in white. The background shows the meshes employed and the normalized contours of the Euclidean norm of the gradient $\nabla J$. We observe that, as expected, both approaches retrieve increasingly accurate solutions as the the number of mesh elements increases. The agglomerated polytopic dG approach with 50 elements is qualitatively comparable, if not better, than the adaptive quadrature approach with 450 elements.
@@
Figure 7: Quantitative comparison (based on the unconstrained test case) between the agglomerated polytopic dG approach and a more classical approach based on triangular elements with adaptive quadrature. The figure on the left depicts the pseudo-time evolution of the difference between the objective value $J$ as shapes are optimized and (an accurate approximation of) the objective value at the optimal solution. The figure on the right depicts the pseudo-time evolution of an approximate distance between the zero-level set of the optimized shapes and the zero-level set of the optimal shape. Markers denote gradient evaluations. We observe that the agglomerated polytopic dG generally converges more rapidly and accurately than the adaptive quadrature approach.